My biggest weakness in math is problems with cylinders so please help me out.
A cylindrical container including a top and bottom is made from no more than 108(pi) cm-squared of sheet metal. If the radius of the top is 3cm what is the maximum volume that the container can hold?

Answer by stanbon(75887)

A cylindrical container including a top and bottom is made from no more than 108(pi) cm-squared of sheet metal. If the radius of the top is 3cm what is the maximum volume that the container can hold?
------------------------
If radius is 3 cm, area of the top is 9pi sq. cm.
Top + bottom = 18pi sq.cm.
=============================
Surface area of the side is 108pi cm^2 - 18pi cm^2 = 90pi cm^2
------------------------------
Formula for the surface of the area = 2pi*r*h
So, 2pi*r*h cm^2 = 90pi cm^2
rh = 45
But r=3
So, h = 45/3 = 15 cm
--------------------------
Volume = pir^2h
Volume = pi*3^2*15
Volume = 135pi cm^3
===================
Cheers,
Stan H.